# Locus: Agentic Predicate Synthesis for Directed Fuzzing

### v2

@@ 4.5. RQ4: Detecting New Vulnerabilities
We integrate Locus into a real-world vulnerability detection workflow and apply the pipeline to a set of well-fuzzed targets, such as VLC(VideoLAN, 2025), libarchive(libarchive contributors, 2025), and libming(libming contributors, 2025). To construct meaningful fuzzing targets, we generate new canaries through two primary approaches. First, we leverage alerts produced by the static analysis tool SVF(Sui and Xue, 2016), which identify potentially vulnerable program points based on memory access patterns, aliasing behavior, or use-after-free risks. Second, we derive reachability canaries for program points associated with previously patched bugs. The rationale behind this strategy is that many real-world bugs occur in clusters or evolve from incomplete fixes(Wang et al., 2020).
We launch 42 fuzzing samples in total, with each fuzzing campaign lasting for 24 hours. We found seven previously undiscovered bugs, including memory leaks, use-after-free, and out-of-bound access (Table 7). At the time of writing, we responsibly reported all the new vulnerabilities to the relevant maintainers, and one of the maintainers has already drafted a fix patch. To further evaluate the efficiency improvement, we also repeat the fuzzing campaign by fuzzers without Locus. The results show that without Locus, AFL++ alone can only detect 2/7 vulnerabilities, and SelectFuzz alone can only detect 3/7 vulnerabilities. Section 4.6 elaborates on one newly detected vulnerability.
Table 7. New vulnerabilities detected fuzzers with Locus. For each newly found vulnerability, we run AFL++ and SelectFuzz w/ and w/o Locus. We exclude the detailed software information per vulnerability report guidelines.
ID Vul. Type AFL++ SelectFuzz
@@ 4.6. Case Study
Figure 3. Locus can sometimes synthesize canary conditions even more precisely than those in Magma, even when provided only with the security patch. Prior to the fix, a dequoting flag was erroneously set to 1, enabling attackers to access uninitialized memory by forcing dequoting.
@@
Canary generation by LocusSection 4.5 discusses our attempts to generate canaries for arbitrary types of vulnerability definitions without relying on existing ones. This shows that Locus does not rely on a pre-defined canary to be applicable. Instead, it can be extended to automatically generate canaries based on diverse representations of the target states, such as the patch and bug descriptions.
@@
To apply Locus in such settings, we also consider letting Locus generate canary conditions based on available security patches. Specifically, we iterate over all security patches in Magma and prompt Locus to automate the generation of a canary condition (where Magma relied on manual analysis), such that the unpatched version satisfies the generated canary, while the patched version violates it. By manually checking all the canaries generated by the LLMs, we confirm that 27 out of 28 security patches can be correctly translated into the same condition.
@@
We find the remaining case to be intriguing: the LLM produces a more precise vulnerability canary than the original one in the Magma benchmark, as shown in Figure 3). Specifically, the vulnerability is found in SQLite3 and classified as CWE-908, i.e., use of the uninitialized resource. When building a collate token, a constant value 1 is incorrectly passed in instead of the dequote flag. This allows attackers to access uninitialized memory via forcibly dequoting, which would lead to memory corruption. The original vulnerability canary only checks whether dequoting is disabled, and invalidates cases for unquoted names since they would not be dequoted even if the dequote flag were set.
Instead, the canary generated by Locus focuses specifically on cases in which a quoted string is mishandled. Such a canary characterizes the real vulnerability more precisely and can potentially reduce the false positives introduced by the imprecise ground truth canaries.
@@
Real-world vulnerability  We elaborate on the vulnerability libarchive-fm54 in Table 7, an out-of-bounds error can occur when the program processes a specially crafted RAR file. The vulnerability stems from insufficient bound checking in the delta filter logic used during decompression. In the RAR format, filters are lightweight data transformation routines applied to improve compression efficiency. One common example is the delta filter, which is often used on binary data, such as audio or image streams, to convert absolute values into relative differences, thus making patterns more compressible. During decompression, the filter operates on two memory regions: the source buffer src, which holds the raw decompressed data, and the destination buffer dst, which stores the filtered output. This vulnerability will be triggered when src inadvertently points into the dst buffer. This overlap can lead to undefined behavior, including memory corruption, out-of-bounds access, or program crashes, as the filter may read from or write to memory locations outside the valid range.
@@
Figure 4 shows the predicates synthesized by Locus that check the input is a valid RAR format file and that the delta filter processing logic will be triggered during decompression. It is seemingly more helpful to generate the predicate to compare the address of src and dst. However, these pointers are only assigned and computed within the vulnerable function itself. By the time this function is reached, it can be too late for such a predicate to effectively guide the fuzzer.
@@ 5. Related Work
Direct grey-box fuzzing  Directed grey-box fuzzing is particularly challenging, primarily due to the prohibitively large search space with sparse rewards(Li et al., 2024c; Luo et al., 2023), i.e., unlike undirected fuzzing where any new coverage indicates progress. Most prior works develop heuristics through static analysis by computing path distances to target states to narrow down the search space (She et al., 2024; Shah et al., 2022; Huang et al., 2024a; Böhme et al., 2017; Luo et al., 2023; Srivastava et al., 2022b; Li et al., 2024c; Chen and Chen, 2018; Lee et al., 2021). For example, AFLGo(Böhme et al., 2017) introduces the idea of using distance metrics between test inputs and target basic blocks to prioritize seeds that are closer to the target, while CAFL(Lee et al., 2021) further improves this metric by introducing specialized progress-capturing state representation and calculating the distance from the testing inputs to the nearest state.
@@
LLM-based fuzzing  LLMs have demonstrated promising code reasoning capabilities(Ding et al., 2024c; Wang et al., 2024; Ding et al., 2024b; Gong et al., 2024; Gu et al., 2024; Yang et al., 2025b). There has been growing interest in applying them to support fuzzing (DARPA, 2024). Most existing approaches utilize LLMs to directly generate test inputs for the target program (Zhang et al., 2024c; Hu et al., 2023; Zhang et al., 2024a; Yang et al., 2023, 2024a; Deng et al., 2024, 2023; Xia et al., 2024) or build grammar-aware input generators (i.e., fuzzing harness) to constrain the search space (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d). While these methods demonstrate the potential of LLMs to accelerate fuzzing, they often require LLMs to reason about the target program states all the way to the inputs, making them susceptible to fundamental challenges in program analysis, such as path explosions and alias dependencies (Jiang et al., 2024; Ding et al., 2024a; Li et al., 2024a; Risse and Böhme, 2024), while also pose challenges to LLMs in terms of the bloated context length and unbounded errors.
@@
LLM-driven proof synthesisLocus shares a similar spirit with recent research that focuses on reasoning about loop invariants and program specifications to synthesize proofs and automate program verification(Chakraborty et al., 2023; Chen et al., 2024; Lu et al., 2024; Kozyrev et al., 2024; Yang et al., 2024c). Locus resembles these approaches in the sense that they also integrate LLMs with rule-based verifiers, such as theorem provers or SMT solvers, to check the validity of synthesized properties. However, Locus relaxes the requirement that the synthesized specifications must facilitate rigorous proof steps, but treats the predicates as best-effort guidance for input search. It bridges the gap between formal verification and dynamic testing, leveraging the reasoning capabilities of LLMs to identify meaningful program states while prioritizing efficiency and generality over formal guarantees.
@@ 6. Threats to Validity
Threats to LLM types  Second, our evaluation is currently limited to a specific set of LLMs and benchmark programs. The generality of our results to other models, especially open-weight LLMs and broader software systems, remains to be validated. The performance of our approach may depend on the underlying LLM capability and training data. Additionally, our evaluated benchmark Magma, though built on popular software and real-world vulnerabilities, may not capture the comprehensive challenges present in diverse software.
@@ 7. Conclusion
This paper presented Locus, a novel framework for enhancing directed fuzzing through synthetic progress-capturing predicates. With an agentic synthesizer-validator architecture, Locus effectively guides fuzzing via the predicates while also ensuring any errors in the synthetic predicates are fixed by symbolic execution. Our evaluation demonstrated that Locus substantially improves the state-of-the-art fuzzers. So far, it has uncovered seven previously unpatched vulnerabilities across three software, with one already acknowledged with a draft patch.

### v3

@@ 4.5. RQ4: Detecting New Vulnerabilities
We integrate Locus into a real-world vulnerability detection workflow and apply the pipeline to a set of well-fuzzed targets, such as VLC(VideoLAN, 2025), libming(libming contributors, 2025), libarchive(libarchive contributors, 2025), and tcpreplay (Turner and Contributors, 2001). To construct meaningful fuzzing targets, we generate new canaries through two primary approaches. First, we leverage alerts produced by the static analysis tool SVF(Sui and Xue, 2016), which identify potentially vulnerable program points based on memory access patterns, aliasing behavior, or use-after-free risks. Second, we derive reachability canaries for program points associated with previously patched bugs. The rationale behind this strategy is that many real-world bugs occur in clusters or evolve from incomplete fixes(Wang et al., 2020).
We launch 48 fuzzing samples in total, with each fuzzing campaign lasting for 24 hours. We found nine previously undiscovered bugs, including memory leaks, use-after-free, null pointer dereference, and out-of-bound memory access (Table 6). At the time of writing, we responsibly reported all the new vulnerabilities to the relevant maintainers, and three of the bugs already have drafted fixes. To further evaluate the efficiency improvement, we also repeat the fuzzing campaign by fuzzers without Locus. The results show that without Locus, AFL++ alone can only detect 2/9 vulnerabilities, and SelectFuzz alone can only detect 5/9 vulnerabilities. Section 4.6 elaborates on one newly detected vulnerability.
Table 6. New vulnerabilities detected by fuzzers with Locus. For each newly found vulnerability, we run AFL++ and SelectFuzz with and without Locus.
Bug ID Type AFL++ SelectFuzz
@@
tcpreplay-pmgq NULL deref T.O. 14219 T.O. T.O.
tcpreplay-pg55 OOB access T.O. 31526 20392 21305
@@ 4.6. Case Study
Canary generation by LocusSection 4.5 discusses our attempts to generate canaries for arbitrary target states without relying on existing ones. This shows that Locus need not rely on a pre-defined canary to be applicable. Instead, it can be extended to automatically generate canaries based on diverse representations of the target states, such as patch and bug descriptions.
Figure 3. Locus generates a more precise canary using only the security patch. Previously, an incorrectly set dequoting flag enabled access to uninitialized memory.
@@
To apply Locus in such settings, we also consider letting Locus generate canary conditions based on available security patches. Specifically, we iterate over all security patches in Magma and prompt Locus to automate the generation of a canary condition (Magma relies on manual analysis), such that the unpatched version satisfies the generated canary, while the patched version violates it. By manually checking all the canaries generated by the LLMs, we confirm that 27 out of 28 security patches can be correctly translated into the same condition.
@@
We find the remaining case to be intriguing: the LLM produces a more precise vulnerability canary than the original one in the Magma benchmark, as shown in Figure 3). Specifically, this vulnerability is found in SQLite3 and classified as CWE-908, i.e., use of the uninitialized resource. When building a collate token, a constant value 1 is incorrectly passed in instead of the dequote flag. This allows attackers to access uninitialized memory via forced dequoting, which would lead to memory corruption. The original vulnerability canary only checks whether dequoting is disabled, and invalidates cases for unquoted names since they would not be dequoted even if the dequote flag were set. Instead, the canary generated by Locus targets mishandled quoted strings, precisely capturing the underlying vulnerability and potentially reducing false positives from imprecise ground-truth canaries.
@@
Real-world vulnerability  In this section, we elaborate on the vulnerability libarchive-fm54 in Table 6. An out-of-bounds error can occur when the program processes a specially crafted RAR file. This vulnerability stems from insufficient bound checking in the delta filter logic used during decompression. In the RAR format, filters are lightweight data transformation routines applied to improve compression efficiency. One common example is the delta filter, which is often used on binary data, such as audio or image streams, to convert absolute values into relative differences, thus making patterns more compressible. During decompression, the filter operates on two memory regions: the source buffer src, which holds the raw decompressed data, and the destination buffer dst, which stores the filtered output. This vulnerability will be triggered when src inadvertently points into the dst buffer. This overlap can lead to undefined behavior, including memory corruption, out-of-bounds access, or program crashes, as the filter may read from or write to memory locations outside the valid range.
@@
Figure 4 shows the predicates synthesized by Locus that check the input is a valid RAR format file and that the delta filter processing logic will be triggered during decompression. A predicate comparing the addresses of src and dst seems more helpful, but these pointers are set and computed inside the vulnerable function. By the time execution reaches it, the predicate is too late to guide the fuzzer effectively.
@@
Apply Locus’s design to other languagesLocus analyzes program behaviors using LLMs, which have been shown to have strong performance on code analysis tasks across various programming languages (Jimenez et al., 2024; Ding et al., 2024b; Gu et al., 2024). In this section, we include preliminary results on extending Locus to broader programming languages. We replace the synthesizer agent with Claude Code (Anthropic, 2024) to ensure greater compatibility. Since KLEE offers limited support for non-C languages, we employ a self-reflection strategy that enables the agent to validate the semantic correctness of generated predicates.
Due to the limited availability of directed fuzzers for other languages, we apply Locus to three coverage-guided fuzzers in other ecosystems: Go-fuzz (Vyukov, 2015) for Go, Jazzer (GmbH and Contributors, 2020) for Java, and Atheris (Google and Contributors, 2020) for Python. For each language, we select two publicly reported vulnerabilities and use LLMs to generate canaries from the corresponding fixing patches (consistent with Section 4.6). We use Claude Sonnet 4.5 as the agent model.
[…]
Table 7. TTE for applying Locus to other languages
Vulnerability Language Baseline Locus
idna#108 Python 14 5
CVE-2020-35655 Python T.O. 36509
hjson#17 Go 18354 2298
gojay#32 Go T.O. 74732
CVE-2021-44228 Java T.O. T.O.
CVE-2021-37714 Java T.O. T.O.
@@ 5. Related Work
Directed grey-box fuzzing  Directed grey-box fuzzing is challenging, primarily due to the prohibitively large search space with sparse rewards(Li et al., 2024c; Luo et al., 2023), i.e., unlike coverage-guided fuzzing where any new coverage indicates progress. Most prior works develop heuristics through static analysis by computing distances to target states to narrow down the search space (She et al., 2024; Shah et al., 2022; Huang et al., 2024a; Böhme et al., 2017; Luo et al., 2023; Srivastava et al., 2022b; Li et al., 2024c; Chen and Chen, 2018; Lee et al., 2021). For example, AFLGo(Böhme et al., 2017) uses distance metrics between test inputs and target basic blocks to prioritize seeds that are closer to the target, while CAFL(Lee et al., 2021) further improves this metric by introducing specialized progress-capturing state representation and calculating the distance from the testing inputs to the nearest state.
@@
LLM-based fuzzing  LLMs have demonstrated promising code understanding and analysis capabilities(Ding et al., 2024d; Wang et al., 2024; Ding et al., 2024c; Gong et al., 2024; Gu et al., 2024; Yang et al., 2025b). There has been growing interest in applying them to support fuzzing (DARPA, 2024). Most existing approaches utilize LLMs to directly generate test inputs for the target program (Zhang et al., 2024a; Hu et al., 2023; Zhang et al., 2024b; Yang et al., 2023, 2024a; Deng et al., 2024, 2023; Xia et al., 2024) or build grammar-aware input generators (i.e., fuzzing harness) to constrain the search space (Liu et al., 2024a; Huang et al., 2024b; Liu et al., 2024c; Shi et al., 2024; Zhang et al., 2024d). While these methods demonstrate the potential of LLMs to accelerate fuzzing, they often require LLMs to reason about the target program states all the way to the inputs, making them susceptible to fundamental challenges in program analysis, such as path explosions and alias dependencies (Jiang et al., 2024; Ding et al., 2024a; Li et al., 2024a; Risse and Böhme, 2024), while also pose challenges to LLMs in terms of the bloated context length and unbounded errors.
@@
LLM-driven proof synthesisLocus shares a similar spirit with recent research that focuses on reasoning about loop invariants and program specifications to synthesize proofs and automate program verification(Chakraborty et al., 2023; Chen et al., 2024; Lu et al., 2024; Kozyrev et al., 2024). Locus resembles these approaches in the sense that they also integrate LLMs with rule-based verifiers, such as theorem provers or SMT solvers, to check the validity of synthesized properties. However, Locus relaxes the requirement that the synthesized specifications must facilitate rigorous proof steps, but treats the predicates as best-effort guidance for input search. It bridges the gap between formal verification and dynamic testing, leveraging the reasoning capabilities of LLMs to identify meaningful program states while prioritizing efficiency and generality over formal guarantees.
@@ 6. Threats to Validity
Threats to LLM types  Second, our evaluation covers only a limited set of LLMs and benchmark programs. The generality of our results to other models, especially open-weight LLMs and broader software systems, remains to be validated. The performance of our approach may depend on the underlying LLM capability and training data. Additionally, our evaluated benchmark Magma, though built on popular software and real-world vulnerabilities, may not capture the comprehensive challenges present in diverse software.
@@ 7. Conclusion
This paper presented Locus, a novel framework for enhancing directed fuzzing through synthetic progress-capturing predicates. With an agentic synthesizer-validator architecture, Locus effectively guides fuzzing via the predicates while also ensuring any errors in the synthesized predicates are fixed by symbolic execution. Our evaluation demonstrated that Locus substantially improves the state-of-the-art fuzzers. So far, it has uncovered nine previously unpatched vulnerabilities across three software projects, with three already acknowledged with draft fixes.
Acknowledgements.
We are grateful to Jun Yang, Weichen Li, Chenghao Yang, Chenyuan Yang, Zheng Yu, Heqing Huang, Penghui Li, Weiteng Chen, and Miltos Allamanis for sharing their constructive and insightful feedback, which significantly helped improve this paper. This research is supported in part by Open Philanthropy, NSF CAREER Award CNS-2442719, and generous gifts from OpenAI. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors. Results presented in this paper were obtained using the Chameleon testbed supported by the National Science Foundation. Language model assistants were used to help polish the paper.